# High Dimensional Ensemble Kalman Filter

### v1

@@ 4.2 Convergence of High Dimensional Ensemble Kalman Filter
To show the impact of the dimensionality on the MSE, we take the banding and the mid-banding estimators as an example. If $q\asymp p$, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O((n^{-1}{\log p})^{2\alpha/(2(\alpha+1))})$, which implies that the MSE between the analysis states of the EnKF $\hat{\mathbf{x}}_{t}^{a}$ and the $\mathbf{x}_{t}^{a}$ mainly depends on the estimation error of the Kalman gain matrix. Commonly in high resolution data assimilation, the dimension of the observation vector $\mathbf{y}_{t}$ is of much smaller order than that of the state vector $\mathbf{x}_{t}$, so that $q\ll p$. Assume $q=p^{\beta}$ for a $\beta\in(0,1)$, then the one-step assimilation MSE can be written as
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}n^{-1}+n^{-1}p^{\beta-1}{n}^{1/(\alpha+1)}(\log p)^{\alpha/(\alpha+1)}\big{)}.$
If $\beta$ is large so that $p^{\beta-1}{n}^{1/(\alpha+1)}(\log p)^{\alpha/(\alpha+1)}\rightarrow\infty$, the estimation error of the Kalman gain matrix will dominate the MSE, that is, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\big{)}$. If $\beta$ is small such that $p^{\beta-1}{n}^{1/(\alpha+1)}(\log p)^{\alpha/(\alpha+1)}\rightarrow 0$, then $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O(n^{-1})$. For the impact of the observation dimension $q$ on the MSE, a study reported in the SM shows that if the additional observation $y_{t,q+1}$ is not far away from its forecast ${\mathbf{H}}_{t}\mathbf{x}_{t}^{f}$, the MSE will decrease as the first term in (4.1) will decrease as $q$ increases.
@@ 5 Imperfect Models
Theorem 5.1 presents the bounds of the difference between the forecast/analysis states of the EnKF and the Kalman Filter for the one-step assimilation when $\mathcal{M}_{t}$ is unknown. Denote
@@
Suppose Assumptions 1(i) or 1(ii), 2, 3, 5 hold and the analysis ensemble $\{\mathbf{x}_{t-1,j}^{a}\}_{j=1}^{n}$ is generated independently from the distribution of $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}$ with mean $\mathbf{x}_{t-1}^{a}$ and covariance $\boldsymbol{\Sigma}_{t-1}^{a}$ at time $t-1$. Then, the forecast and analysis states of the EnKF satisfy
@@
[…]
$\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}n^{-1}+\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}(\|\Xi_{\xi_{t}}\|^{2}+n^{-1})\big{)}=O\big{(}n^{-1}+\|\Xi_{\xi_{t}}\|^{2})\big{)}.$ (5.4)
Thus, if $\|\Xi_{\xi_{t}}\|^{2}=O(n^{-1})$, then $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}=O\left(n^{-1}\right)=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$, which means that the model error does not alter the order of magnitude of the assimilation and forecasting MSEs of the EnKF. However, if $n\|\Xi_{\xi_{t}}\|^{2}\rightarrow\infty$, then $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O(\|\Xi_{\xi_{t}}\|^{2})$, which means the modelling error dominates the upper bounds of the MSE in the analysis states of the EnKF. When the dimensions $p$ and $q\rightarrow\infty$,
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}\asymp\max\{n^{-1},\;p^{-1}{q}\tilde{\Delta}_{\mathbf{K},t},\;p^{-1}{q}\|\Xi_{\xi_{t}}\|^{2}\}.$ (5.5)
[…]
@@
The above analysis shows that for both the fixed and the diverging dimension, and both the unbiased and biased approximation of $\mathcal{M}_{t}$, the approximation errors as represented by $\mu_{\xi_{t}}$ and $\|\Xi_{\xi_{t}}\|^{2}$ contributes substantially to the MSE of the assimilated $\hat{\mathbf{x}}_{t}^{a}$ by the EnKF, which highlights the benefits of having more accurate model operators. While this may be obvious, Theorem 5.1 describes exactly how it happens. In summary, the approximation error of the model operator $\mathcal{M}_{t}$ has a strong influence on the MSE of the data assimilation of the EnKF. This suggests that a good approximation of the model operator is a significant aspect of data assimilation. We also introduce an iterative EnKF update algorithm in S.1.2 to account for the model approximation error to some extent since most geophysical models are imperfect in practice.
@@ 6 Simulation
This section reports results of simulation studies performed on two popular geophysical models, the Lorenz-96 and the Shallow Water Equation models, to demonstrate the proposed high dimensional EnKF algorithms. These two models are often used for testing new approaches in atmospheric sciences for both data assimilation and the spatial-temporal forecasts. The proposed high dimensional EnKF methods are compared with the standard EnKF and the inflation method (Liang et al., 2012) under both the correct and imperfect models. Details of the proposed high dimensional EnKF algorithms are given in Section S.1 of the SM, including the consistent EnKF update in S.1.1 and the iterative EnKF update algorithms in S.1.2, which are based on the consistent estimators of the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding estimators.
@@ 6.1 Lorenz-96 model
Lorenz-96 model (L96) (Lorenz, 1996) is a continuous in time but discrete in space model for modeling spatially extended chaotic systems in the atmosphere. It is a nonlinear model for continuous evolution of atmospheric variables such as temperature or vorticity at $p$ locations on a latitude circle defined by the ordinary differential equation
${d\mathbf{x}_{t}^{(j)}}/{dt}=(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}-\mathbf{x}_{t}^{(j)}+F\overset{\triangle}{=}f_{j}(\mathbf{x}_{t},t),$ (6.1)
@@
for $j=1,\ldots,p$, with $\mathbf{x}_{t}^{(-1)}=\mathbf{x}_{t}^{(p-1)},\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}^{(p)}$ and $\mathbf{x}_{t}^{(p+1)}=\mathbf{x}_{t}^{(1)}$, where $\mathbf{x}_{t}^{(j)}$ denote the $j^{th}$ dimension of the state variable $\mathbf{x}_{t}$, $(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}$ is the nonlinear term, $-\mathbf{x}_{t}^{(j)}$ is the damping term and $F$ is the force term. The $F$ can be set as other constants to generate imperfect models, the context discussed in Section 5 .
@@
L96 can be solved approximately via the fourth-order Runge-Kutta time integration scheme (RK4; Butcher, 2016), which approximates the model as $\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)=\mathbf{x}_{t}+h(k_{1}+2k_{2}+2k_{3}+k_{4})/{6},$ where $k_{1}=\mathbf{f}\left(\mathbf{x}_{t},t\right)$, $k_{2}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{1}/2,t+h/2\right)$, $k_{3}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{2}/2,t+h/2\right)$, $k_{4}=\mathbf{f}\left(\mathbf{x}_{t}+hk_{3},t+h\right)$ and $\mathbf{f}=\left(f_{1},f_{2},\dots,f_{p}\right)^{\mathrm{T}}$. An approximation to the true state was simulated by either (i) implementing the RK4 method with a finer time step, say $h=0.05$, which corresponds to $6$ hours in the real-time assuming that the time-scale of the dissipation in the atmosphere is $5$ days (Lorenz, 1996); or (ii) by adding a random noise at each time step so that $\mathbf{x}_{t+1}=\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)+\mathbf{w}_{t+1}$. In our experiment, $\mathbf{w}_{t}\sim\mathrm{N}(0,\sigma_{0}\mathbf{I}_{p})$ and $\sigma_{0}$ were set as $0.1,0.25$ and $0.5$, respectively.
@@
Furthermore, we experimented $p=40$ and $100$, under which the system behaves chaotically, since the L96 model behaves like chaos for $p\geq 7$ (Bedrossian, Blumenthal and Punshon-Smith, 2022). The true forcing was $F=8$. Similar to Liang et al. (2012), we also considered the misspecified model setting by making $F=6,7,9,10$ in the assimilation processes, corresponding to the imperfect models in Section 5. The initial state $\mathbf{x}_{1}$ was such that $x_{1}^{(j)}=F$ for $j\neq 20$ and $x_{1}^{(20)}=F+0.001$, and the initial forecast ensemble was generated by $\mathbf{x}_{1,j}^{f}=\mathbf{x}_{1}+\mathrm{N}_{p}(0,0.1\mathbf{I}_{p})$ for $j=1,\ldots,n$.
@@
The observations $\mathbf{y}_{t}$ were made at $q$ randomly selected components of the state vector $\mathbf{x}_{t}$ with $q=30,40$ and $100$ while $p=40$ and $100$, respectively, followed by adding observation errors $\varepsilon_{t}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{q}(0,\mathbf{R}_{t})$ to the true states. The $(i,j)-$element of $\mathbf{R}_{t}$ was $R_{ij}=0.5^{\min\left\{|i-j|,p-|i-j|\right\}}$, which prescribes spatially correlated observation errors to suit geophysical data. The observation operator $\mathbf{H}_{t}$ was a binary $q\times p$ matrix with the $(j,i)$ element being $1$ if and only if the $j-$th component of $\mathbf{y}_{t}$ was the observed $i-$th dimension of the state $\mathbf{x}_{t}$. We simulated observations every $4$ time steps for a total of $2000$ steps and reported the results for the last $1000$ steps to allow the system to settle down. The ensemble size $n$ was $20,30,60$ and $90$. All the simulation experiments were repeated 500 times.
@@
We considered the high dimensional EnKF schemes using the banding, mid-banding, tapering and thresholding covariance estimators, which were compared with the standard EnKF that utilizes the sample covariance to estimate the forecast error covariance without the proposed high dimensional treatment and the inflation methods (Liang et al., 2012). The performances of the assimilation methods were evaluated by the root-mean-square error (RMSE) of the analysis state $\hat{\mathbf{x}}_{t}^{a}$ to (i) the underlying analysis state $\mathbf{x}_{t}^{a}$ by the Kalman Filter utilizing the true covariances and (ii) the true state $\mathbf{x}_{t}$, averaged over the last 1000 time steps.
However, the underlying analysis states $\mathbf{x}_{t}^{a}$ were unknown in the simulation of L96 due to its nonlinear nature. Indeed, for a nonlinear $\mathcal{M}_{t}$, the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ could not be readily updated by (2.3) in the Kalman Filter. To gain an accurate estimate of $\boldsymbol{\Sigma}_{t}^{f}$ for the oracle analysis $\mathbf{x}_{t}^{a}$, we used the standard EnKF with a very large ensemble size $1000$ (relative to $p$ in the simulation), and used the sample covariance matrices to replace $\boldsymbol{\Sigma}_{t}^{f}$ in the Kalman filter update formula to obtain $\mathbf{x}_{t}^{a}$, since the large ensemble size would make the sample covariance a high quality substitute for $\boldsymbol{\Sigma}_{t}^{f}$.
@@
Figure 2 displays the average RMSEs to the oracle analysis state of the five assimilation methods, namely the standard EnKF, the inflation method (Liang et al., 2012) and the three high dimensional EnKF equipped with the banding and tapering (both respect the circular bandable structures), and the thresholding estimators, with respect to different forcing level $F$ used in the model at moderate dimension $p=40$ and high dimension $p=100$ under different $\sigma_{0}$ and $n$. It shows that the EnKF with the three high dimensional covariance estimators had much smaller RMSEs than the standard EnKF and the inflation methods at both the correct ($F=8$) and misspecified ($F\neq 8$) model settings. Even when $F$ became more deviated from the underlying value $8$, the analysis states obtained by the high dimensional EnKFs were still much closer to the oracle analysis states than the inflation method although both strains of methods showed increased RMSEs with respect to $|F-8|$. The standard EnKF with the sample covariance estimator to $\boldsymbol{\Sigma}_{t}^{f}$ had much larger assimilation error in all cases.
@@
Comparing Figure 2’s panels (d)-(f) with panels (a)-(c), we see that the advantages of the proposed high dimensional EnKF schemes over the inflation method were more apparent for the higher dimension setting $p=100$. The analysis RMSEs of the high dimensional EnKF methods kept at a lower level especially when the force $F$ was equal or close to the true value ($F=8$), while the analysis RMSE of the inflation method ranked behind but better than the standard EnKF. This means that the proposed high dimensional EnKF methods are quite attractive and useful as the geophysical models in earth science are of ultra high dimension. The figure also displays the percentages of divergence of the five methods, which shows that the inflation method endured quite high divergences, even higher than the standard EnKF. In contrast, the three high dimensional EnKF methods had quite low divergence rates for all cases. The divergence rate of the inflation method increases substantially along with the increase of the model error, which, for instance, reached 50% when $\sigma=0.5$.

### v2

@@ 3.2 Convergence of High Dimensional Ensemble Kalman Filter
To show the impact of the dimensionality on the MSE, we take the banding and mid-banding estimators as examples. If $q\asymp p$, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O((n^{-1}{\log p})^{2\alpha/(2(\alpha+1))})$, which implies that the one-step assimilation MSE mainly depends on the estimation error of the Kalman gain matrix. Commonly in high resolution data assimilation, the dimension of the observation $\mathbf{y}_{t}$ is much smaller than that of the state vector $\mathbf{x}_{t}$, so that $q\ll p$. If $q$ is relatively large such that $p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\}\gg n^{-1}$, then the estimation error of the Kalman gain matrix will dominate, that is, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\big{)}$. Otherwise, the first term $n^{-1}$ dominates. For the impact of the observation dimension $q$ on the MSE, a study reported in the SM shows that if the additional observation $y_{t,q+1}$ is not far away from its forecast ${\mathbf{H}}_{t}\mathbf{x}_{t}^{f}$, the MSE will decrease as the first term in (3.1) will decrease as $q$ increases.
For the discrepancies between the EnKF analysis state and the true state, as $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}+\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ and $\mathbf{x}_{t}$ is a random variable with $\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{a})$ which is a constant not related to the ensemble size $n$ for a fixed $p$, then $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ can not converge to zero.
In practice, the data assimilation process can be repeated for multiple time steps ahead to make longer-time predictions. The following Theorem 3.2 gives the $h$-step MSE bounds between the forecast/analysis states of the EnKF in comparison to the underlying Kalman Filters. To simplify writing, define two non-stochastic constants
$C_{t,h}^{A}=\frac{1-L_{t,h}^{h}A_{t,h}^{h}}{1-L_{t,h}A_{t,h}}\,\text{ and }C_{t,h}^{B}=\frac{1-L_{t,h}^{h}B_{t,h}^{h}}{1-L_{t,h}B_{t,h}},$ (3.7)
where $A_{t,h}=\max_{t<k\leq t+h}\|\mathbf{I}-\mathbf{K}_{k}\mathbf{H}_{k}\|$, $B_{t,h}=\max_{t<k\leq t+h}\{\mathbb{E}\|\mathbf{I}-\hat{\mathbf{K}}_{k}\mathbf{H}_{k}\|\}$, and $L_{t,h}=\max_{0<k\leq h}\{\ell_{0},\ell_{t+k}\}$ for a fixed $t$, with $\ell_{0}$ and $\ell_{t}$ defined in Assumption 1. We further denote
$U_{k}=\big{\{}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\Delta_{\mathbf{K},k}+\operatorname{tr}\big{(}\mathbb{E}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})^{\mathrm{T}}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})({\mathbf{H}}_{k}\boldsymbol{\Sigma}_{k}^{f}{\mathbf{H}}_{k}^{\mathrm{T}}+\mathbf{R}_{k})\big{)}\big{\}}^{1/2}.$ (3.8)
Theorem 3.2.
Assumptions 1(i) or 1(ii), 2, 3 hold and the analysis ensemble $\{\mathbf{x}_{t,j}^{a}\}_{j=1}^{n}$ is generated independently from the distribution of $\mathbf{x}_{t}\mid\mathbf{y}_{1:t}$ with mean $\mathbf{x}_{t}^{a}$ and covariance $\boldsymbol{\Sigma}_{t}^{a}$ at time $t$. By repeatedly performing the ensemble forecast-analysis process $h$ steps, the forecast and analysis states of the EnKF at the time $t+h$ satisfy
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{f}-\mathbf{x}_{t+h}^{f})|\mathbf{y}_{1:t+h-1}\|^{2}\leq p^{-1}\Big{(}\{\operatorname{tr}(n^{-1}\boldsymbol{\Sigma}_{t+h}^{f})\}^{1/2}+L_{t,h}C_{t,h-1}^{B}\max_{t<k\leq t+h-1}U_{k}\Big{)}^{2}$
$=O\Big{(}n^{-1}+p^{-1}L_{t,h}^{2}(C_{t,h-1}^{A})^{2}\max_{t<k\leq t+h-1}\big{\{}\Delta_{\mathbf{K},k}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\big{\}}\Big{)},$
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}\leq p^{-1}\Big{(}\{\operatorname{tr}(n^{-1}\boldsymbol{\Sigma}_{t+h}^{a})\}^{1/2}+C_{t,h}^{B}\max_{t<k\leq t+h}U_{k}\Big{)}^{2}$
$=O\Big{(}n^{-1}+p^{-1}(C_{t,h}^{A})^{2}\max_{t<k\leq t+h}\big{\{}\Delta_{\mathbf{K},k}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\big{\}}\Big{)}.$
Compared with the one-step results given in Theorem 3.1, the term related to the estimation error of $\hat{\boldsymbol{\Sigma}}_{t+k}^{f}$ in the $h$-step analysis MSE is inflated with constants $(C_{t,h}^{A})^{2}$ which are related to $L_{t,h}$ and $A_{t,h}$ defined before (3.7). However, for any fixed $h$, $\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{a}-\mathbf{x}_{t+h}^{a})|\mathbf{y}_{1:t+h}\|^{2}\asymp\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}.$
Sometimes, the assimilation process proceeds every $k$ time steps not every step as one may not have observations at each step or the assimilation in every step is too computationally consuming. Theorem S.3.2 of the SM gives the MSE bounds for multiple-step forecasting and assimilation where the assimilation process proceeds every $k$ time steps.
@@ 4 Imperfect Models
Theorem 4.1 presents the one-step assimilation bounds of the difference between the forecast/analysis states of the EnKF and the Kalman Filter when $\mathcal{M}_{t}$ is unknown. Denote
@@
Suppose Assumptions 1(i) or 1(ii), 2, 3, 4 hold and $\{\mathbf{x}_{t-1,j}^{a}\}_{j=1}^{n}$ is generated independently from the distribution of $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}$ with mean $\mathbf{x}_{t-1}^{a}$ and covariance $\boldsymbol{\Sigma}_{t-1}^{a}$ at time $t-1$. Then, the forecast and analysis states of the EnKF satisfy
@@
The theorem suggests that the model approximation error creates extra terms related to $\mu_{\xi_{t}}$ and $\|\boldsymbol{\Xi}_{\xi_{t}}\|$, which are absent from the results in Theorem 3.1 where accurate model is assumed. Specifically, the assimilation MSEs for fixed $p,q$ and $p,q\rightarrow\infty$ are, respectively,
@@
[…]
The above analysis shows that for both the fixed and diverging dimensions and both the unbiased and biased approximation of $\mathcal{M}_{t}$, the approximation errors as represented by $\mu_{\xi_{t}}$ and $\|\Xi_{\xi_{t}}\|^{2}$ contributes substantially to the assimilation MSE, which suggests that a good approximation of the model operator is a significant aspect of data assimilation. While this may be obvious, Theorem 4.1 describes exactly how it happens. The accumulation of the model approximation error is more pronounced in the $h-$step assimilation MSEs as shown in Theorem S.3.3 of the SM. Thus, an iterative HD-EnKF algorithm is introduced in S.1.2 to account for the model approximation error to some extent since most geophysical models are imperfect in practice, of which the main idea is to substitute $\hat{\mathbf{x}}_{t}^{f}$ in the original forecast sample covariance matrix by the analysis state $\hat{\mathbf{x}}_{t}^{a}$ in the iteration.
@@ 5 Simulation
This section reports results of simulation studies performed on two popular geophysical models, the Lorenz-96 and the Shallow Water Equation models, to demonstrate the proposed HD-EnKF algorithms. These two models are often used for testing new approaches in atmospheric sciences for both data assimilation and the spatial-temporal forecasts. The proposed HD-EnKF methods are compared with the standard EnKF and the inflation method (Liang et al., 2012) under both the correct and imperfect models. Details of the proposed HD-EnKF algorithms are given in Section S.1 of the SM, including the consistent HD-EnKF update in S.1.1 and the iterative HD-EnKF update algorithms in S.1.2, which are based on the consistent estimators of the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the banding, mid-banding, tapering and thresholding estimators.
@@ 5.1 Lorenz-96 model
Lorenz-96 model (L96; Lorenz, 1996) is a continuous in time but discrete in space model for modeling spatially extended chaotic systems in the atmosphere. It is a nonlinear model for continuous evolution of atmospheric variables such as temperature or vorticity at $p$ locations on a latitude circle defined by the ordinary differential equation
${d\mathbf{x}_{t}^{(j)}}/{dt}=(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}-\mathbf{x}_{t}^{(j)}+F\overset{\triangle}{=}f_{j}(\mathbf{x}_{t},t),\,\text{ for }j=1,\ldots,p,$ (5.1)
@@
with $\mathbf{x}_{t}^{(-1)}=\mathbf{x}_{t}^{(p-1)},\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}^{(p)}$ and $\mathbf{x}_{t}^{(p+1)}=\mathbf{x}_{t}^{(1)}$, where $\mathbf{x}_{t}^{(j)}$ denotes the $j^{th}$ dimension of the state variable $\mathbf{x}_{t}$, $(\mathbf{x}_{t}^{(j+1)}-\mathbf{x}_{t}^{(j-2)})\mathbf{x}_{t}^{(j-1)}$ is the nonlinear term, $-\mathbf{x}_{t}^{(j)}$ is the damping term and $F$ is a forcing term.
@@
L96 can be solved approximately via the fourth-order Runge-Kutta time integration scheme (RK4; Butcher, 2016). An approximation to the true state was simulated by (i) implementing the RK4 method with a finer time step, say $h=0.05$; and (ii) adding a random noise at each time step so that $\mathbf{x}_{t+1}=\mathcal{M}_{t+1}\left(\mathbf{x}_{t}\right)+\mathbf{w}_{t+1}$. In our experiment, $\mathbf{w}_{t}\sim\mathrm{N}(0,\sigma_{0}\mathbf{I}_{p})$ and $\sigma_{0}$ were set as $0.1,0.25$ and $0.5$, respectively.
@@
We experimented $p=40$ and $100$, under which the system behaves chaotically, since the L96 model behaves like chaos for $p\geq 7$ (Bedrossian et al., 2022). The true forcing was $F=8$. We also considered the misspecified model setting by making $F=6,7,9,10$ in the assimilation processes, corresponding to the imperfect models in Section 4. The initial state was $x_{1}^{(j)}=F$ for $j\neq 20$ and $x_{1}^{(20)}=F+0.001$, and the initial forecast ensemble was generated by $\mathbf{x}_{1,j}^{f}=\mathbf{x}_{1}+\mathrm{N}_{p}(0,0.1\mathbf{I}_{p})$ for $j=1,\ldots,n$.
@@
The observations $\mathbf{y}_{t}$ were made at $q$ randomly selected components of the state vector $\mathbf{x}_{t}$ with $q=30,40$ and $100$ while $p=40$ and $100$, respectively, followed by adding observation errors $\boldsymbol{\varepsilon}_{t}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{q}(0,\mathbf{R}_{t})$ to the true states. The $(i,j)-$element of $\mathbf{R}_{t}$ was $R_{ij}=0.5^{\min\left\{|i-j|,q-|i-j|\right\}}$, which prescribes spatially correlated observation errors. The observation operator $\mathbf{H}_{t}$ was a $q\times p$ matrix with the $(j,i)$ element being $1$ if and only if the $j-$th component of $\mathbf{y}_{t}$ was the observed $i-$th dimension of the state $\mathbf{x}_{t}$. We simulated observations every $4$ time steps for a total of $2000$ steps and reported the results for the last $1000$ steps to allow the system to settle down. The ensemble size $n$ was $20,30,60$ and $90$. All the experiments were repeated 500 times.
@@
We considered the HD-EnKF schemes using the banding, mid-banding, tapering and thresholding covariance estimators, which were compared with the standard EnKF that utilizes the sample covariance to estimate the forecast error covariance and the inflation methods (Liang et al., 2012). The performances of these methods were evaluated by the root-mean-square error (RMSE) of the analysis state $\hat{\mathbf{x}}_{t}^{a}$ to (i) the underlying analysis state $\mathbf{x}_{t}^{a}$ by the Kalman Filter utilizing the true covariances and (ii) the true state $\mathbf{x}_{t}$. To gain accurate estimations of $\boldsymbol{\Sigma}_{t}^{f}$ for the oracle analysis $\mathbf{x}_{t}^{a}$, we used the standard EnKF with a very large ensemble size $1000$ (relative to $p$ in the simulation), and used the sample covariance matrices to replace $\boldsymbol{\Sigma}_{t}^{f}$ in the Kalman filter update formula to obtain $\mathbf{x}_{t}^{a}$.
@@
Figure 2 displays the average RMSEs to the oracle analysis state of the five assimilation methods, namely the standard EnKF, the inflation method and the three HD-EnKFs equipped with the banding and tapering (both respect the circular bandable structures), and the thresholding estimators, under different forcing $F$ at moderate dimension $p=40$ and high dimension $p=100$ with different $\sigma_{0}$ and $n$. It shows that the HD-EnKF with the three high dimensional covariance estimators had much smaller RMSEs than the standard EnKF and the inflation methods at both the correct ($F=8$) and misspecified ($F\neq 8$) models. Even when $F$ deviated more from the true value $8$, the HD-EnKFs still achieved analysis states that were much closer to the oracle than the inflation method although both strains of methods showed increased RMSEs with respect to $|F-8|$. The standard EnKF with the sample covariance estimator to $\boldsymbol{\Sigma}_{t}^{f}$ had much larger assimilation errors in all cases.
@@
Comparing Figure 2’s panels (d)-(f) with panels (a)-(c), the advantages of the proposed HD-EnKFs over the inflation method were more apparent for the higher dimension $p=100$. The analysis RMSEs of the HD-EnKFs kept at a lower level especially when the force $F$ was equal or close to the true value ($F=8$), while the analysis RMSE of the inflation method ranked behind but better than the standard EnKF. This means that the HD-EnKFs are quite attractive as the geophysical models in earth science are of ultra high dimension. The figure also displays the percentages of divergence of the five methods. It shows that the inflation method endured quite high divergences which increased substantially with the model error (e.g. reached 50% when $\sigma=0.5$), even higher than the standard EnKF. In contrast, the three HD-EnKF methods had quite low divergence rates for all cases.